# On Evidence for Elastic Interactions in the Dark Sector

### v1

@@ 1 Introduction
Traditionally, cosmology has served as fertile ground for theoretical speculation, only limited by the bounds of imagination. This speculative nature of cosmology was largely due to a scarcity of data guiding theoreticians. However, this scenario has undergone a dramatic transformation in recent decades, owing to tremendous technological advancements that have enabled the construction of better instruments for observing the cosmos. The observations gathered with these instruments have delineated the realms of permissible speculation. Paradigmatic examples of how data have shaped our comprehension of the universe include the anisotropies in the Cosmic Microwave background (CMB) by satellites such as WMAP [1] and Planck [2], as well as the mapping of Large Scale Structures (LSS) of the universe through surveys like SDSS [3], DES [4, 5, 6], KiDS [7, 8], HSC [9, 10] or DESI [11]. These observations have not only consolidated $\Lambda$CDM as the standard model of cosmology, but they have also established robust evidence for the existence of a dark sector comprised of dark matter and dark energy.
@@
While cosmological observations have indeed tempered the speculative nature of cosmological models, it is also a common occurrence in physics that an increase in the precision of the data comes with the emergence of anomalies that potentially challenge our physical understanding. Initially, these anomalies may not appear particularly significant, and some may even dissipate with additional data or more refined analysis. However, in other instances, seemingly innocuous anomalies can ultimately pave the way for groundbreaking discoveries, potentially precipitating a paradigm shift. The appearance of some tensions in cosmological data [12, 13], being particularly pressing the $H_{0}$ and the $\sigma_{8}$ tensions, make it plausible that cosmology is presently experiencing one of these revolutionary periods, thus being a unique opportunity to discover novel phenomena. Given the elusive character of the dark sector, it is particularly alluring to seek for new physics associated to dark matter and/or dark energy. This brings us back to theoretical speculation, but this time being guided by very restrictive data.
@@
In this brief review, we will be concerned with the $\sigma_{8}$ tension. As we reach the completion of Stage-III LSS experiments that have provided data with remarkable precision, we witness how inferences from different sources appear to indicate that the clustering of dark matter seems to be systematically weaker than the one expected from extrapolating CMB measurements within $\Lambda$CDM. The amount of clustering can be conveniently encoded in the parameter $\sigma_{8}$ (or its relative $S_{8}$) that gives a measure of the amplitude of matter fluctuations in the present universe and the values reported by low-redshift probes show a trend to be systematically smaller than the value measured by the Planck satellite [12, 13]. These observations might signal that matter clusters less than expected between matter-radiation decoupling and today. Interestingly, the Atacama Cosmology Telescope (ACT) [14] has recently reported values of $\sigma_{8}$ from CMB lensing measurements that are compatible with Planck. The relevance of this measurement is that ACT probes linear scales for redshifts $z\simeq 0.5-5$ with its sensitivity peaking at $z\sim 2$, thus suggesting that, if matter indeed clusters less than expected, the mechanism responsible for such a clustering deficit should be at work in the very late universe $z\lesssim 2$.
@@ 2 A proxy for dark elastic interactions
There are different manners to realise the idea of having elastic interactions in the dark sector. One option is to assume the presence of elastic scatterings between dark matter and dark energy particles [16]. However, the corresponding interaction rate decreases with the expansion so it is only important at high redshift and it dies away as the universe expands. Thus, it does not fully comply with our requirements. More generally, one can study the possible interactions in the dark sector between dark matter characterised as a fluid and dark energy described in terms of a scalar field. This was done in a systematic way in e.g. [17, 18] where the possible interactions were classified into three types, one of which corresponds to pure momentum transfer between the dark components. Different versions of this scenario as well as extensions have been studied in e.g. [19, 20, 21, 22, 23] and similar scenarios where dark energy is described by a vector rather than a scalar field have been considered in [24, 25, 26, 27]. An Effective Field Theory approach was employed in [28] to describe certain interacting scenarios with dark energy and modified gravity. A more phenomenological scenario for elastic interactions at the level of the conservation equations was introduced in [29] and a Lagrangian formulation of similar models was developed in [30]. Barring some particularities of each model, the scenarios featuring an elastic interaction in the dark sector share some common interesting properties. In particular, as anticipated, they can generically take advantage of the dark energy pressure to produce a weaker growth of structures, thus providing compelling scenarios for alleviating the $\sigma_{8}$ tension. It is worth mentioning that models with an elastic coupling between dark energy and baryons have also been explored in e.g. [31, 32]. It was shown in [32] that a reduction of the matter clustering can also take place, although less prominently than in the dark matter case due to the smaller abundance of baryons.
@@
where $\alpha$ is some constant parameter that measures the strength of the interaction and $u^{\mu}_{a}$ is the 4-velocity of the corresponding fluid. For dark matter we will utilise the usual assumptions of being cold, so its pressure and sound speed are nearly vanishing, while for dark energy we will assume a constant equation of state parameter $w$ and some adiabatic sound speed $c_{\text{de}}^{2}$, which we will eventually set to 1 in order to have pressure perturbations able to efficiently prevent the clustering. It is apparent from Eqs. (1) that the interaction only operates when there is a relative motion between the dark components, i.e., on sub-Hubble scales, while the background evolution is not affected by the interaction. The perturbed dark sector described by the density contrasts $\delta_{i}=\delta\rho_{i}/\rho_{i}$ and the peculiar velocities $\theta_{i}$, where $i$ stands for dark matter or dark energy, is governed, in the Newtonian gauge with potential $\Phi$ and in Fourier space, by the equations [^3] From now on, the parameter $\alpha$ will be measured in its natural units $\alpha\to\frac{\alpha}{\rho_{c}H_{0}}$, with $\rho_{c}$ and $H_{0}$ the critical density and Hubble factor today.
$\delta_{\rm dm}^{\prime}$ $=$ $-\theta_{\rm dm}+3\Phi^{\prime}\,,$ (2)
$\delta_{\rm de}^{\prime}$ $=$ $-3\mathcal{H}(c_{\rm de}^{2}-w)\delta_{\rm de}+3(1+w)\Phi^{\prime}-(1+w)\left(1+9\mathcal{H}^{2}\frac{c_{\rm de}^{2}-w}{k^{2}}\right)\theta_{\rm de}\;,$ (3)
$\theta_{\rm dm}^{\prime}$ $=$ $-\mathcal{H}\theta_{\rm dm}+k^{2}\Phi+\Gamma_{\alpha}(\theta_{\rm de}-\theta_{\rm dm})\;,$ (4)
$\theta_{\rm de}^{\prime}$ $=$ $(3c_{\rm de}^{2}-1)\mathcal{H}\theta_{\rm de}+k^{2}\Phi+\frac{k^{2}c_{\rm de}^{2}}{1+w}\delta_{\rm de}-\Gamma_{\alpha}R(\theta_{\rm de}-\theta_{\rm dm})\,,$ (5)
@@
The function $\Gamma_{\alpha}$ describes the interaction rate, while $R$ is the relative abundance of dark matter to dark energy. As advertised, only the Euler equations are modified by the interaction and the relation between the density contrast and the peculiar velocities is the same as in the standard scenario. This feature is important because it allows to infer the density perturbation from the velocity field as in the standard model, despite having a interacting dark sector. In more general interacting scenarios, this is not the case and additional care must be taken when confronting the model predictions to observations. Let us notice the formal resemblance of the Euler equations (4) and (5) with the equations governing the baryon-photon plasma before decoupling (see e.g. [37]) where Thomson scattering generates an analogous term proportional to the relative velocities. [^4] This analogy has motivated to call it sometimes covariantized dark Thomson interaction. We will refer to it as $\alpha$CDM for obvious reasons. There are some important differences however. While Thomson scattering gives rise to an interaction rate that decreases as the universe expands, $\Gamma_{\alpha}$ grows as $\propto a^{4}$ and this ensures that the universe will enter a dark-coupling epoch at late times as it expands, rather than undergoing a decoupling phase. This is a crucial difference of this scenario with the elastic dark scattering considered in [16] which does exit rather than entering a dark coupled epoch. [^5] This is also the reason why no relevant observational effects were found in the baryons-dark energy elastic scattering studied in [31] while important effects were observed in [32].
@@ 3 Mitigating a tension…
The absence of correlations between the interaction parameter $\alpha$ and background quantities acquires especial relevance for the case of $H_{0}$. This is important because many cosmological scenarios exhibit an anti-correlation between both the $H_{0}$ and the $\sigma_{8}$ tensions so that alleviating one of them exacerbates the other. [^6] It has even been suggested that the $\sigma_{8}$ tension might be an $H_{0}$ tension in disguise [39, 38]. One typical example for this effect occurs when allowing for additional relativistic species as an attempt to solve the $H_{0}$ tension. This extra radiation does alleviate to some extent the $H_{0}$ tension, but this requires to go to regions in the parameter space where $\sigma_{8}$ is even larger, thus worsening the $\sigma_{8}$ tension (see left panel in Figure 3). In the presence of the interaction, this is no longer the case since the interaction operates at low redshift to reduce the clustering without affecting the background cosmology, so the alleviation of $H_{0}$ is untied from the value of $\sigma_{8}$, as Figure 3 clearly shows.
@@ 5 Looking ahead: A smoking gun!
$\vec{v}_{\rm dm}^{\prime}+\mathcal{H}\Big{(}1+\Theta\Big{)}\vec{v}_{\rm dm}+\nabla\Phi=0\,,$ (8)
@@
One could complain that the elastic interaction with dark energy is felt only by dark matter, while the baryons that we actually observe in galaxies are oblivious to it. Thus, the question is whether galaxies are good tracers of the dark matter velocity field or rather, they probe the underlying gravitational potential. Intuitively, one can argue that galaxies are mostly formed at sufficiently high redshift when the interaction does not play any role. Once the galaxy has formed and virialises together with the dark matter halo, it will be attached to it. This means that galaxies will follow the dark matter haloes, thus being good tracers of the dark matter velocity field and the interaction is not expected to be strong enough to disrupt the galaxy-halo system. Of course, this intuition should be confirmed by a proper $N$-body simulation. To test this intuition, we have modified the RAMSES code to include the effects of the interaction [44, 45]. In order to distinguish between baryons and dark matter particles, we have implemented two matter species and only for one of them we have included the interaction with dark energy. In Figure 5 we confirm that our intuition is correct and indeed baryons remain attached to the dark matter haloes. As a bonus, in Figure 6 we show how the interacting model predicts a deficit of massive haloes at low redshift. The lack of structures from $N$-body simulations in scenarios with elastic interactions has also been observed in [46, 47, 48].
@@ 6 Discussion
Acknowledgements: It is a pleasure to thank Miguel Aparicio Resco, Wilmar Cardona, Enea Di Dio, Antonio L. Maroto, David F. Mota, Shinji Tsujikawa and Hans Winther for collaborations on these topics. We also thank Carlos García García and Sunny Vagnozzi for useful discussions. This work was supported by the Project PID2021-122938NB-I00 funded by the Spanish “Ministerio de Ciencia e Innovación” and FEDER “A way of making Europe” and the Project SA097P24 funded by Junta de Castilla y León. DB and JBJ also acknowledge support from project PIC-2022-02 funded by Salamanca University. DB also acknowledge support from project PID2022-139841NB-I00 funded by the Spanish “Ministerio de Ciencia e Innovación”. DF acknowledges support from the programme Ayudas para Financiar la Contratación Predoctoral de Personal Investigador (ORDEN EDU/601/2020) funded by Junta de Castilla y León and European Social Fund.

### v2

@@ Abstract
[…]
[5]organization=Centre for Space Research, North-West University, city=Potchefstroom, postcode=2520, country=South Africa
@@ 1 Introduction
Traditionally, cosmology has served as fertile ground for theoretical speculation, only limited by the bounds of imagination. This speculative nature of cosmology was largely due to a scarcity of data guiding theoreticians. However, this scenario has undergone a dramatic transformation in recent decades, owing to tremendous technological advancements that have enabled the construction of better instruments for observing the cosmos. The observations gathered with these instruments have delineated the realms of permissible speculation. Paradigmatic examples of how data have shaped our comprehension of the universe include the anisotropies in the Cosmic Microwave background (CMB) by satellites such as WMAP [1] and Planck [2], as well as the mapping of Large Scale Structures (LSS) of the universe through surveys like SDSS [3], DES [4, 5, 6], KiDS [7, 8], HSC [9, 10] or DESI [11]. These observations have not only helped establishing $\Lambda$CDM as the standard model of cosmology, but they have also provided robust evidence for the existence of a dark sector comprised of dark matter and dark energy.
@@
While cosmological observations have indeed tempered the speculative nature of cosmological models, it is also a common occurrence in physics that an increase in the precision of the data comes with the emergence of anomalies that potentially challenge our physical understanding. Initially, these anomalies may not appear particularly significant, and some may even dissipate with additional data or more refined analysis. However, in other instances, seemingly innocuous anomalies can ultimately pave the way for groundbreaking discoveries, potentially precipitating a paradigm shift. The appearance of some tensions in cosmological data [12, 13], being particularly pressing the $H_{0}$ and the $\sigma_{8}$ tensions although the more recent DESI data has contributed a potential new ”$w_{0}-w_{a}$” tension [14, 15], make it plausible that cosmology is presently experiencing one of these revolutionary periods, thus being a unique opportunity to discover novel phenomena. Given the elusive character of the dark sector, it is particularly alluring to seek for new physics associated to dark matter and/or dark energy. This brings us back to theoretical speculation, but this time being guided by very restrictive data.
@@
In this brief review, we will be concerned with the $\sigma_{8}$ tension. As we reach the completion of Stage-III LSS experiments that have provided data with remarkable precision, we witness how inferences from different sources appear to indicate that the clustering of dark matter seems to be systematically weaker than the one expected from extrapolating CMB measurements within $\Lambda$CDM. The amount of clustering can be conveniently encoded in the parameter $\sigma_{8}$ (or its relative [^1] Although we will refer to the $\sigma_{8}$ and $S_{8}$ tensions as somewhat interchangeable terminology referring to the same cosmological conundrum apparently indicating a clustering deficit, the precise relation between both quantities depends on $\Omega_{\text{m}}$ so, strictly speaking, they are only equivalent if $\Omega_{\text{m}}$ is known. See e.g. [16, 17, 18] for relevant discussions related to the role of $\Omega_{\text{m}}$ in the cosmological tensions. $S_{8}$) that gives a measure of the amplitude of matter fluctuations in the present universe and the values reported by low-redshift probes show a trend to be systematically smaller than the value measured by the Planck satellite [12, 13]. These observations might signal that matter clusters less than expected between matter-radiation decoupling and today. Interestingly, the Atacama Cosmology Telescope (ACT) [19] has recently reported values of $\sigma_{8}$ from CMB lensing measurements that are compatible with Planck. The relevance of this measurement is that ACT probes linear scales for redshifts $z\simeq 0.5-5$ with its sensitivity peaking at $z\sim 2$, thus suggesting that, if matter indeed clusters less than expected, the mechanism responsible for such a clustering deficit should be at work in the very late universe $z\lesssim 2$.
@@ 2 A proxy for dark elastic interactions
There are different manners to realise the idea of having elastic interactions in the dark sector. One option is to assume the presence of elastic scatterings between dark matter and dark energy particles [21]. However, the corresponding interaction rate decreases with the expansion so it is only important at high redshift and it dies away as the universe expands. Thus, it does not fully comply with our requirements. More generally, one can study the possible interactions in the dark sector between dark matter characterised as a fluid and dark energy described in terms of a scalar field. This was done in a systematic way in e.g. [22, 23] where the possible interactions were classified into three types, one of which corresponds to pure momentum transfer between the dark components. Different versions of this scenario as well as extensions have been studied in e.g. [24, 25, 26, 27, 28, 29] and similar scenarios where dark energy is described by a vector rather than a scalar field have been considered in [30, 31, 32, 33]. An Effective Field Theory approach was employed in [34] to describe certain interacting scenarios with dark energy and modified gravity. A more phenomenological scenario for elastic interactions at the level of the conservation equations was introduced in [35] and a Lagrangian formulation of similar models was developed in [36]. Barring some particularities of each model, the scenarios featuring an elastic interaction in the dark sector share some common interesting properties. In particular, as anticipated, they can generically take advantage of the dark energy pressure to produce a weaker growth of structures, thus providing compelling scenarios for alleviating the $\sigma_{8}$ tension. It is worth mentioning that models with an elastic coupling between dark energy and baryons have also been explored in e.g. [37, 38]. It was shown in [38] that a reduction of the matter clustering can also take place, although less prominently than in the dark matter case due to the smaller abundance of baryons. Elastic interactions of dark matter with photons [39, 40, 41] and neutrinos [42] have also been explored. However, in all these cases the interactions are more tightly constrained because they give rise to novel scattering channels that would affect e.g. measurements from collider experiments.
@@
where $\alpha$ is some constant parameter that measures the strength of the interaction and $u^{\mu}_{a}$ is the 4-velocity of the corresponding fluid. For dark matter we will utilise the usual assumptions of being cold, so its pressure and sound speed are nearly vanishing, while for dark energy we will assume a constant equation of state parameter $w$ and some sound speed $c_{\text{de}}^{2}$, which we will eventually set to 1 in order to have pressure perturbations able to efficiently prevent the clustering. It is apparent from Eqs. (1) that the interaction only operates when there is a relative motion between the dark components, i.e., on sub-Hubble scales, while the background evolution is not affected by the interaction. The perturbed dark sector described by the density contrasts $\delta_{j}=\delta\rho_{j}/\rho_{j}$ and the peculiar velocities $\theta_{j}=i\vec{k}\cdot\vec{v}_{j}$, where $j$ stands for dark matter or dark energy, is governed, in Fourier space and assuming no anisotropic stresses, by the equations [^4] From now on, the parameter $\alpha$ will be measured in its natural units $\alpha\to\frac{\alpha}{\rho_{c}H_{0}}$, with $\rho_{c}$ and $H_{0}$ the critical density and Hubble factor today.
$\delta_{\rm dm}^{\prime}$ $=$ $-\theta_{\rm dm}+h_{A}\,,$ (2)
$\delta_{\rm de}^{\prime}$ $=$ $-3\mathcal{H}\left(c_{\rm de}^{2}-w\right)\delta_{\rm de}+(1+w)h_{A}-(1+w)\left(1+9\mathcal{H}^{2}\frac{c_{\rm de}^{2}-w}{k^{2}}\right)\theta_{\rm de}\;,$ (3)
$\theta_{\rm dm}^{\prime}$ $=$ $-\mathcal{H}\theta_{\rm dm}+k^{2}\Phi+\Gamma_{\alpha}\Big{(}\theta_{\rm de}-\theta_{\rm dm}\Big{)}\;,$ (4)
$\theta_{\rm de}^{\prime}$ $=$ $\left(3c_{\rm de}^{2}-1\right)\mathcal{H}\theta_{\rm de}+k^{2}\Phi+\frac{c_{\rm de}^{2}k^{2}}{1+w}\delta_{\rm de}-\Gamma_{\alpha}R\Big{(}\theta_{\rm de}-\theta_{\rm dm}\Big{)}\,,$ (5)
@@
The above perturbation equations are valid in both Newtonian and synchronous gauge with the metric perturbation $h_{A}$ taking the values, in the notation of [47], $h_{\text{Newtonian}}=3\Phi^{\prime}$ and $h_{\text{synchronous}}=-\frac{1}{2}h^{\prime}$ and the understanding that the term $k^{2}\Phi$ in (4) and (5) should be removed in the synchronous gauge. Our discussions on the interacting scenario below will be mainly referred to the Newtonian gauge. However, an important property of the interaction is its gauge invariance (at linear order) because velocity perturbations change with the time derivative of the spatial gauge parameter, so the gauge dependence disappears when we take the difference of any two velocity perturbations. In other words, although we need a reference frame to define a velocity, the relative velocities do not depend on the frame choice. Thus, the interaction term can be defined and analysed in any frame and it is not affected by possible superfluous gauge modes as those present in e.g. the synchronous gauge. The function $\Gamma_{\alpha}$ describes the interaction rate, while $R$ is the relative abundance of dark matter to dark energy. As advertised, only the Euler equations are modified by the interaction and the relation between the density contrast and the peculiar velocities is the same as in the standard scenario. This feature is important because it allows to infer the density perturbation from the velocity field as in the standard model, despite having an interacting dark sector. In more general interacting scenarios, this is not the case and additional care must be taken when confronting the model predictions to observations. Let us notice the formal resemblance of the Euler equations (4) and (5) with the equations governing the baryon-photon plasma before decoupling (see e.g. [48]) where Thomson scattering generates an analogous term proportional to the relative velocities. [^5] This analogy has motivated to call it sometimes covariantized dark Thomson interaction. We will refer to it as $\alpha$CDM for obvious reasons. There are some important differences however. While Thomson scattering gives rise to an interaction rate that decreases as the universe expands, $\Gamma_{\alpha}$ grows as $\propto a^{4}$ and this ensures that the universe will enter a dark-coupling epoch at late times as it expands, rather than undergoing a decoupling phase. This is a crucial difference of this scenario with the elastic dark scattering considered in [21] which does exit rather than entering a dark coupled epoch. [^6] This is also the reason why no relevant observational effects were found in the baryons-dark energy elastic scattering studied in [37] while important effects were observed in [38]. Some observational constraints on this model have been obtained in [49].
@@ 3 Mitigating a tension…
The absence of correlations between the interaction parameter $\alpha$ and background quantities acquires especial relevance for the case of $H_{0}$. This is important because many cosmological scenarios exhibit an anti-correlation between both the $H_{0}$ and the $\sigma_{8}$ tensions so that alleviating one of them exacerbates the other (see e.g. [51, 17, 18]). [^7] It has even been suggested that the $\sigma_{8}$ tension might be an $H_{0}$ tension in disguise [52, 50]. One typical example for this effect occurs when allowing for additional relativistic species as an attempt to solve the $H_{0}$ tension. This extra radiation does alleviate to some extent the $H_{0}$ tension, but this requires to go to regions in the parameter space where $\sigma_{8}$ is even larger, thus worsening the $\sigma_{8}$ tension (see left panel in Figure 3). In the presence of the interaction, this is no longer the case since the interaction operates at low redshift to reduce the clustering without affecting the background cosmology, so the alleviation of $H_{0}$ is untied from the value of $\sigma_{8}$, as Figure 3 clearly shows.
@@ 5 Looking ahead: A smoking gun!
$\vec{v}^{\>\prime}_{\rm dm}+\mathcal{H}\Big{(}1+\Theta\Big{)}\vec{v}_{\rm dm}+\nabla\Phi=0\,,$ (8)
@@
One could complain that the elastic interaction with dark energy is felt only by dark matter, while the baryons that we actually observe in galaxies are oblivious to it. Thus, the question is whether galaxies are good tracers of the dark matter velocity field or rather, they probe the underlying gravitational potential. Intuitively, one can argue that galaxies are mostly formed at sufficiently high redshift when the interaction does not play any role. Once the galaxy has formed and virialises together with the dark matter halo, it will be attached to it. This means that galaxies will follow the dark matter haloes, thus being good tracers of the dark matter velocity field and the interaction is not expected to be strong enough to disrupt the galaxy-halo system. Of course, this intuition should be confirmed by a proper $N$-body simulation. To test this intuition, we have modified the RAMSES code to include the effects of the interaction [57, 58]. In order to distinguish between baryons and dark matter particles, we have implemented two matter species and only for one of them we have included the interaction with dark energy. In Figure 5 we confirm that our intuition is correct and indeed baryons remain attached to the dark matter haloes. As a bonus, in Figure 6 we show how the interacting model predicts a deficit of massive haloes at low redshift. The lack of structures from $N$-body simulations in scenarios with elastic interactions has also been observed in [59, 60, 61, 62].
@@ 6 Discussion
Acknowledgements: It is a pleasure to thank Miguel Aparicio Resco, Wilmar Cardona, Enea Di Dio, Antonio L. Maroto, David F. Mota, Shinji Tsujikawa and Hans Winther for collaborations on these topics. We also thank Pierre Fleury, Carlos García García and Sunny Vagnozzi for useful discussions. This work was supported by the Project PID2021-122938NB-I00 funded by the Spanish “Ministerio de Ciencia e Innovación” and FEDER “A way of making Europe” and the Project SA097P24 funded by Junta de Castilla y León. DB and JBJ also acknowledge support from project PIC-2022-02 funded by Salamanca University. DB also acknowledge support from project PID2022-139841NB-I00 funded by the Spanish “Ministerio de Ciencia e Innovación”. DF acknowledges support from the programme Ayudas para Financiar la Contratación Predoctoral de Personal Investigador (ORDEN EDU/601/2020) funded by Junta de Castilla y León and European Social Fund and the financial support provided by the postdoctoral research fellowship NWU PDRF Fund NW.1G01487 by the North-West University and the Centre of Space Research
Codes availability: Modified versions of CAMB and CLASS including the elastic interactions are available upon request.